With a head wind, a plane traveled 1000 miles in 4 hours. With the same wind as a tail wind, the return trip took 3 hours and 20 minutes, Find the plane's air speed and the wind speed.
Answer by Alan3354(69443)

1000/4 = 250 mph
1000/(10/3) = 300 mph
The plane's airspeed is the average (250 + 300)/2 = 275 mph
The windspeed is the difference, 275 - 250 = 25 mph
-----------------------
headwind is one word.
so is tailwind.